How to solve the quadratic equation for x by factoring?
For Instance, 6x^2 + 3x = 0

Answer by CharlesG2(834)

How to solve the quadratic equation for x by factoring?
For Instance, 6x^2 + 3x = 0

6x^2 + 3x = 0
divide both sides by 6
x^2 + (1/2)x = 0
we need to complete the square
take 1/2 the coefficient on the x term, square it and add to both sides
1/2 * 1/2 = 1/4, (1/4)^2 = 1/16
which gives us:
x^2 + (1/2)x + 1/16 = 1/16
which factors to:
(x + 1/4)^2 = (+- 1/4)^2, by FOIL this is the same as the line above
x + 1/4 = 1/4 or x + 1/4 = -1/4
x = 0 or x = -1/2